# Graph Data Condensation via Self-expressive Graph Structure Reconstruction

### v1

@@ Abstract.
With the increasing demands of training graph neural networks (GNNs) on large-scale graphs, graph data condensation has emerged as a critical technique to relieve the storage and time costs during the training phase. It aims to condense the original large-scale graph to a much smaller synthetic graph while preserving the essential information necessary for efficiently training a downstream GNN. However, existing methods concentrate either on optimizing node features exclusively or endeavor to independently learn node features and the graph structure generator. They could not explicitly leverage the information of the original graph structure and failed to construct an interpretable graph structure for the synthetic dataset. To address these issues, we introduce a novel framework named Graph Data Condensation via Self-expressive Graph Structure Reconstruction (GCSR). Our method stands out by (1) explicitly incorporating the original graph structure into the condensing process and (2) capturing the nuanced interdependencies between the condensed nodes by reconstructing an interpretable self-expressive graph structure. Extensive experiments and comprehensive analysis validate the efficacy of the proposed method across diverse GNN models and datasets. Our code is available at https://www.dropbox.com/scl/fi/2aonyp5ln5gisdqtjimu8/GCSR.zip?rlkey=11cuwfpsf54wxiiktu0klud0x&dl=0.
Graph Data Condensation, Graph Neural Network, Node Classification
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: SIGKDD Conference on Knowledge Discovery and Data Mining; June 2 2023; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data mining
@@ 1. Introduction
In recent years, the significant increase in the volume of graph data in different fields, such as social networks (Naseem et al., 2021; Lamsal, 2021), recommendation systems (Wu et al., 2020; Eirinaki et al., 2018), traffic networks (Liang et al., 2023; Liu et al., 2023), and knowledge graphs (Zhang et al., 2020; Wang et al., 2018a), has presented challenges for practitioners and researchers. Consequently, storing, processing, analyzing, and transmitting large-scale data has become a burden in real-world applications. Moreover, when considering deep learning tasks such as neural architecture search (Ren et al., 2021) and continual learning (Han et al., 2020; Li and Hoiem, 2017), directly utilizing the full dataset for training a graph neural network (GNN) has the potential to result in poor efficiency.
@@ 3. Preliminary
Graph Condensation. A graph dataset $\mathcal{T}$ is denoted as $\mathcal{T}=\{\textbf{A},\textbf{X},\textbf{Y}\}$, where $\textbf{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix, $\textbf{X}\in\mathbb{R}^{N\times d}$ is the node feature, and $\textbf{Y}\in\mathbb{R}^{N\times 1}$ is the label. Generally, graph data condensation aims to learn a downsized synthetic graph $\mathcal{S}=\{\textbf{A}^{\prime},\textbf{X}^{\prime},\textbf{Y}^{\prime}\}$ with $\textbf{A}^{\prime}\in\mathbb{R}^{N^{\prime}\times N^{\prime}}$, $\textbf{X}^{\prime}\in\mathbb{R}^{N^{\prime}\times d}$, and $\textbf{Y}^{\prime}\in\mathbb{R}^{N^{\prime}\times 1}$ from the original training graph $\mathcal{T}$ ($N^{\prime}\ll N$). Both $\textbf{Y}_{i}$ and $\textbf{Y}^{\prime}_{j}$ belong to the class set $\mathcal{C}=\{1,2,\cdots,C\}$. Formally, graph condensation is defined by solving the problem as follows:
@@ 4.1.1. Node Initialization
Moreover, to make the synthetic dataset conform to the distribution of the original dataset and avoid creating an unbalanced dataset, we keep the class distribution the same as the original graph. Formally, the following equation holds for every node class, i.e., $c\in\{1,2,\cdots,C\}$.
@@ 4.1.2. Regularizer Initialization
15:  end for// Computing the final adjacency matrix $\mathbf{A}^{\prime}$
@@ 5.2. Overall Performance
For dataset condensation baselines, all the synthetic graphs are generated by SGC. We reuse the most of results in (Jin et al., 2021) since the experiment setting is exactly the same. We re-implement SFGC and SGDD to guarantee a fair comparison since they use GCN to generate the condensed graph dataset. The comparison of all the baselines is shown in Table 2. From the table, we can observe that: (1) overall, GCSR achieves superior performance compared to the baselines. The condensed data is comparable to the original dataset for training a GNN for node classification. (2) Compared to baselines, GCSR exhibits an improvement in accuracy with an increase in the condensed graph size, which contradicts the observations reported in prior studies and adheres to the intuition that more data leads to better performance. This further validates that our method is effective and successfully solves the graph dataset condensation task. (3) Moreover, compared with baselines that solely update synthetic node features such as DCG, SFGC, and MTT, our approach outperforms them by a large margin. We attribute the improved performance to the effectively generated adjacency matrix that captures both the self-expressiveness property of node features and the information of the original graph structure.
@@ 5.5. Learned Node Feature Visualization
To evaluate whether our method can capture more information from the original graph, we use t-SNE (Van der Maaten and Hinton, 2008) to visualize the real and synthetic node features of Cora condensed by GCond, SFGC, SGDD, and GCSR. For a clearer display, we only show nodes from 3 out of 7 classes. We also calculate the average silhouette coefficient (Rousseeuw, 1987)
(a metric that measures the distance between samples of different classes and a higher coefficient is better) of synthetic data generated from each method. Fig. 5 shows that: (1) data learned by GCond, SFGC, and SGDD is not well conformed to the distribution of the original data, and some nodes are mixed up with nodes of other classes. (2) Data learned by GCSR obeys the distribution of the original data and exhibits a distinct separation between different classes. This could lead to better classification performance and evident distribution characteristics. (3) GCSR achieves the highest silhouette coefficient, indicating the learned node features of different classes are easy to separate.

### v2

@@ Abstract.
With the increasing demands of training graph neural networks (GNNs) on large-scale graphs, graph data condensation has emerged as a critical technique to relieve the storage and time costs during the training phase. It aims to condense the original large-scale graph to a much smaller synthetic graph while preserving the essential information necessary for efficiently training a downstream GNN. However, existing methods concentrate either on optimizing node features exclusively or endeavor to independently learn node features and the graph structure generator. They could not explicitly leverage the information of the original graph structure and failed to construct an interpretable graph structure for the synthetic dataset. To address these issues, we introduce a novel framework named Graph Data Condensation via Self-expressive Graph Structure Reconstruction (GCSR). Our method stands out by (1) explicitly incorporating the original graph structure into the condensing process and (2) capturing the nuanced interdependencies between the condensed nodes by reconstructing an interpretable self-expressive graph structure. Extensive experiments and comprehensive analysis validate the efficacy of the proposed method across diverse GNN models and datasets. Our code is available at https://github.com/zclzcl0223/GCSR.
Graph Data Condensation; Graph Neural Network; Node Classification
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671710isbn: 979-8-4007-0490-1/24/08ccs: Information systems Data mining
@@ 1. Introduction
In recent years, the significant increase in the volume of graph data in different fields, such as social networks (Naseem et al., 2021; Lamsal, 2021), recommendation systems (Wu et al., 2020; Eirinaki et al., 2018), traffic networks (Liang et al., 2023; Liu et al., 2023b, 2024b, a), and knowledge graphs (Zhang et al., 2020; Wang et al., 2018a), has presented challenges for practitioners and researchers. Consequently, storing, processing, analyzing, and transmitting large-scale data has become a burden in real-world applications. Moreover, when considering deep learning tasks such as neural architecture search (Ren et al., 2021; Liu et al., 2024a) and continual learning (Han et al., 2020; Li and Hoiem, 2017), directly utilizing the full dataset for training a graph neural network (GNN) has the potential to result in poor efficiency.
@@ 3. Preliminary
Graph Condensation. A graph dataset $\mathcal{T}$ is denoted as $\mathcal{T}=\{\textbf{A},\textbf{X},\textbf{Y}\}$, where $\textbf{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix, $\textbf{X}\in\mathbb{R}^{N\times d}$ is the node feature, and $\textbf{Y}\in\mathbb{R}^{N\times 1}$ is the label. Generally, graph data condensation aims to learn a downsized synthetic graph $\mathcal{S}=\{\textbf{A}^{\prime},\textbf{X}^{\prime},\textbf{Y}^{\prime}\}$ with $\textbf{A}^{\prime}\in\mathbb{R}^{N^{\prime}\times N^{\prime}}$, $\textbf{X}^{\prime}\in\mathbb{R}^{N^{\prime}\times d}$, and $\textbf{Y}^{\prime}\in\mathbb{R}^{N^{\prime}\times 1}$ from the original training graph $\mathcal{T}$ ($N^{\prime}\ll N$). Both $\textbf{Y}_{i}$ and $\textbf{Y}^{\prime}_{j}$ belong to the class set $\mathcal{C}=\{0,1,\cdots,C-1\}$. Formally, graph condensation is defined by solving the problem as follows:
@@ 4.1.1. Node Initialization
Moreover, to make the synthetic dataset conform to the distribution of the original dataset and avoid creating an unbalanced dataset, we keep the class distribution the same as the original graph. Formally, the following equation holds for every node class, i.e., $c\in\{0,1,\cdots,C-1\}$.
@@ 4.1.2. Regularizer Initialization
1115:  end for
// Computing the final adjacency matrix $\mathbf{A}^{\prime}$
@@ 5.2. Overall Performance
For dataset condensation baselines, all the synthetic graphs are generated by SGC. We reuse most of the results in (Jin et al., 2021) since the experiment setting is exactly the same. We re-implement SFGC and SGDD to guarantee a fair comparison since they use GCN to generate the condensed graph dataset. The comparison of all the baselines is shown in Table 2. From the table, we can observe that: (1) overall, GCSR achieves superior performance compared to the baselines. The condensed data is comparable to the original dataset for training a GNN for node classification. (2) Compared to baselines, GCSR exhibits an improvement in accuracy with an increase in the condensed graph size, which contradicts the observations reported in prior studies and adheres to the intuition that more data leads to better performance. This further validates that our method is effective and successfully solves the graph dataset condensation task. (3) Moreover, compared with baselines that solely update synthetic node features such as DCG, SFGC, and MTT, our approach outperforms them by a large margin. We attribute the improved performance to the effectively generated adjacency matrix that captures both the self-expressiveness property of node features and the information of the original graph structure.
@@ 5.5. Learned Node Feature Visualization
To evaluate whether our method can capture more information from the original graph, we use t-SNE (Van der Maaten and Hinton, 2008) to visualize the real and synthetic node features of Cora condensed by GCond, SFGC, SGDD, and GCSR. For a clearer display, we only show nodes from 3 out of 7 classes. We also calculate the average silhouette coefficient (Rousseeuw, 1987) (a metric that measures the distance between samples of different classes and a higher coefficient is better) of synthetic data generated from each method. Fig. 5 shows that: (1) data learned by GCond, SFGC, and SGDD is not well conformed to the distribution of the original data, and some nodes are mixed up with nodes of other classes. (2) Data learned by GCSR obeys the distribution of the original data and exhibits a distinct separation between different classes. This could lead to better classification performance and evident distribution characteristics. (3) GCSR achieves the highest silhouette coefficient, indicating the learned node features of different classes are easy to separate.
@@ 6. Conclusion
Acknowledgement
This work was sponsored by National Key Research and Development Program of China under Grant No.2022YFB3904204, National Natural Science Foundation of China under Grant No. 62102246, 62272301, and Provincial Key Research and Development Program of Zhejiang under Grant No. 2021C01034. Part of the work was done when the students were doing internships at Yunqi Academy of Engineering.
Appendix A The Effectiveness of Message Passing Initialization
Table 6. The effectiveness of MPI.
Dataset GCSR SFGC SFGC+MPI GCond GCond+MPI
Citeseer (r=1.8%) 74.0 $\pm$ 0.4 70.8 $\pm$ 0.4 71.8 $\pm$ 0.4 $\uparrow$ 69.8 $\pm$ 1.4 71.3 $\pm$ 0.9 $\uparrow$
Reddit (r=0.1%) 91.2 $\pm$ 0.2 84.6 $\pm$ 1.6 90.5 $\pm$ 0.2 $\uparrow$ 89.6 $\pm$ 0.7 90.2 $\pm$ 0.2 $\uparrow$
Table 7. The sensitivity analysis of $k$ in MPI.
Dataset 0 2 4 6 8 10
Citeseer(r=1.8%) 69.2 $\pm$ 1.2 70.2 $\pm$ 0.8 70.2 $\pm$ 0.8 70.3 $\pm$ 0.7 70.4 $\pm$ 0.7 70.3 $\pm$ 0.7
Table 8. The effectiveness of graph regularizer initialization of GCSR.
Method Citeseer (r=1.8%) Reddit (r=0.1%)
GCSR 71.8 $\pm$ 0.9 91.2 $\pm$ 0.2
Sub-structure Init 69.5 $\pm$ 2.0 91.0 $\pm$ 0.3
Random Init 64.8 $\pm$ 2.6 72.2 $\pm$ 2.5
Table 9. The sensitive analysis of regularizer update parameter $\tau$.
Dataset 0.5 0.9 0.99 0.999 1
Citeseer(r=1.8%) 70.2 $\pm$ 1.0 71.8 $\pm$ 0.9 70.2 $\pm$ 1.2 68.7 $\pm$ 1.7 68.1 $\pm$ 1.8
Table 10. The sensitive analysis of regularizer update parameter $\gamma$.
Dataset 0.1 0.5 0.9 1
Citeseer(r=1.8%) 71.6 $\pm$ 0.9 71.8 $\pm$ 0.9 71.6 $\pm$ 0.9 70.6 $\pm$ 0.8
Table 11. The sensitivity analysis of $\mathbf{\alpha}$.
Dataset 0 0.1 0.5 1 10
Citeseer(r=1.8%) 70.1 $\pm$ 0.8 70.2 $\pm$ 1.0 71.1 $\pm$ 0.8 71.8 $\pm$ 0.9 70.1 $\pm$ 1.0
Ogbn-arxiv (r=0.05%) 59.9 $\pm$ 1.1 60.6 $\pm$ 1.1 59.2 $\pm$ 1.3 58.5 $\pm$ 1.8 58.8 $\pm$ 0.7
Table 12. The sensitivity analysis of $\mathbf{\beta}$.
Dataset 0 0.1 0.5 1 10
Citeseer(r=1.8%) 71.6 $\pm$ 0.9 71.8 $\pm$ 0.9 71.6 $\pm$ 1.0 71.6 $\pm$ 0.9 71.3 $\pm$ 1.0
Ogbn-arxiv (r=0.05%) 59.8 $\pm$ 1.3 60.3 $\pm$ 1.3 60.2 $\pm$ 1.3 60.6 $\pm$ 1.1 60.5 $\pm$ 1.0
GCSR introduces message-passing initialization (MPI), which is a novel approach in the field of graph condensation. To further verify the effectiveness of MPI, we implement it on two other graph condensation frameworks and check whether MPI brings performance enhancement. The result is shown in Table 6. We could observe that with MPI, the performance of SFGC and GCond both improve significantly, which validates the effectiveness of our proposed MPI module. Nevertheless, our method still performs better, which illustrates the necessity of SER.
Furthermore, we evaluate the sensitivity of parameter $k$ in the MPI module. We solely implement MPI on GCSR and get the result of different $k$ on Table 7. Consistent with our paper, MPI could significantly improve condensation performance. However, the results are not significantly influenced by $k$. As illustrated in the table above, even if we set $k$ to 8 or 10, the improvement compared to $k$=2 is very small.
Appendix B Graph Regularizer Update Analysis
We initialize the graph regularizer with a probabilistic adjacency matrix to avoid the inconsistency between the dimensions of adjacency matrices of the synthetic graph and the original graph. To further verify the effectiveness of our initialization, we compare our initialization with sub-structure initialization of DosCond (Jin et al., 2022) and random initialization. The result is shown in Table 8. We can observe that the probabilistic adjacency matrix initialization achieves the best performance.
Furthermore, update momentum ratio $\tau$ and $\gamma$ play important roles in updating the graph regularizer in Eqs. 17 and 18. We evaluate their impact by setting different $\tau$ and $\gamma$ as shown in Table 9 and Table 10. Here $\tau=1$ means not update P and $\gamma=1$ means not update $\textbf{Z}_{h}$. We could observe that these two regularizer updates contribute to the final performance. Moreover, the performance is more sensitive to $\tau$, so it is recommended to search for more $\tau$ in the real application.
Appendix C Graph Reconstruction Analysis
In this section, we analyze the impact of hyper-parameter $\alpha$ and $\beta$ in reconstructing the graph structure $\mathbf{Z}$ in Eq. 13 on Citeseer and Ogbn-arxiv as shown in Table 11 and Table 12. Here, $\alpha=0$ means we only apply regularizer $\textbf{Z}_{h}$ and $\beta=0$ indicates we only apply regularizer P. We could observe that adding each regularizer could enhance the performance and the joint use of two regularizers brings the best performance. Moreover, the hyperparameter is recommended to search in (0,1].